Operation Summary for March 3, 2010 - March 5, 2010 |
Wednesday March 3 The day shift began with Operations monitoring the Tevatron (TeV) store 7647 (78.95E30) and the Recycler stash (356.5E10), with the Antiproton Source stacking (15.6E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7647 at 10:06 AM. A Linac RF station (LRF3) tripped off due to a bad Thyratron tube; its repair took 45 minutes. Operations reported multiple MI RF station trips; RF technicians resolved the problem. Operations established store 7649 at 1:08 PM with an Average Initial Luminosity (AIL) of 328.7E30. Operations established beam to Meson MTest (T953) at 2:52 PM. During the evening shift, Operations disabled the MTest beam at 6 PM. LRF5 tripped off a couple of times due to vacuum pressure; it reset. Thursday March 4 The midnight shift began with Operations monitoring store 7649 (94.89E30) and the stash (303.3E10), with Pbar stacking (20.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to MTest at 4 AM. At 5:07 AM, the Linac Klystron Debuncher water skid developed a LCW leak and tripped off. Linac personnel began repairs, which continued through to the end of shift. CDF reported making repairs on a tracking chamber power supply. During the day shift, the Klystron Debuncher water skid repairs continued. Booster personnel accessed for RF repairs. NuMI personnel accessed their target hall for preventative maintenance on the target air system. Linac personnel completed their repairs and Operations had beam to users by 12:49 PM. Operations terminated store 7649 at 2:15 PM and established store 7650 at 3:39 PM with an AIL of 344.4E30. During the evening shift, Operations disabled the MTest beam at 6 PM. Several MI and Linac RF stations tripped off and caused a small amount of downtime; they reset. Friday March 5 The midnight shift began with Operations monitoring store 7650 (116.66E30) and the stash (261.3E10), with Pbar stacking (5.7E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, BRF4 started sparking and needed to be turned off. Column one in CUB went down causing the MI LCW to go into stand-alone mode. Operations established beam to MTest at 4:05 AM. Operations terminated store 7650 at 6:57 AM. Operations put the TeV into shot setup. The Plans for Friday and the Weekend The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/08/2010 email Fermilab |
| Security, Privacy, Legal | |